OpenStudy (anonymous):

Solve the quadratic equation x2 + 12x + 20 = 0 by completing the square.

OpenStudy (superdavesuper):

Completing the square means re-arranging the eqn so that (x+a)^2=b^2 Since (x+a)^2=x^2+2ax+a^2 a=6 in this case....can u go on from here?

OpenStudy (anonymous):

NO I cant

OpenStudy (superdavesuper):

ok i will do one more steps here :) rearranging the eqn x^2 + 12x + 20 = 0 x^2 + 2(6)x + 6^2 -6^2 +20 = 0 u should be able to solve it from here!

OpenStudy (anonymous):

(x^ 2 + 12x) = -20

OpenStudy (anonymous):

Thanks c;

OpenStudy (anonymous):

then divide 12x by 2 and square it should get x^2 +12x +36 = -20 +36

OpenStudy (anonymous):

so (x-6)^2 = 16 so answer should work out to be x = 10 ,2